OpenStudy (anonymous):

Use the transformations of (x)=x^2 to graph the following functions g(x)=4(x-3)^2+3

OpenStudy (callisto):

If you have to graph y=(x-3)^2 using f(x)=x^2, what will you do?

OpenStudy (anonymous):

I have no idea.... need help!!

OpenStudy (callisto):

Test the value of y's by substituting the value of x into the formula For y = x^2 x | -2 | -1 | 0 | 1 | 2 | y | 4 | 1 | 0 | 1 | 4 | For y = (x-3)^2 x | -1 | 0 | 1 | 2 | 3 | 4 | 5 | y | | | | | | | | Can you fill in the row of y for the second table?

OpenStudy (anonymous):

y|1|0|1|1|3|2|5

OpenStudy (anonymous):

is this right......

OpenStudy (callisto):

No o_o! When x=-1, y = [(-1)-3]^2 = (-4)^2 = 16 Can you try it again?

OpenStudy (anonymous):

never mind... to frustrated thanx!!!!

OpenStudy (anonymous):

I just need to see how you suppose to graph the function..... someone please help!!!

OpenStudy (callisto):

f(x-k) => shift f(x) to the right by k units kf(x) => magnify f(x) by k units f(x) + k => shift f(x) upwards by k units. These are the rules you need.

OpenStudy (anonymous):

Okay thanx!!!